# thanks: Corresponding author

### v2

@@ VIII Conclusions
In conclusion, we have proposed to use the “spin-opstring”, extracted from Stochastic Series Expansion QMC as a input data type for machine learning applications. The spin-opstring is a compact representation of the QMC simulation cell that retains all quantum information of the system while being memory-efficient. It consists of the initial state of the simulation cell followed by an operator string that describes the state’s evolution through imaginary time, ultimately bringing it back to its original configuration. In order to demonstrate the applicability of spin-opstring as ML input, we have applied supervised machine learning on two different models : the two-dimensional $t_{2}-V_{1}$ model and the one-dimensional Su-Schrieffer-Heeger (SSH) model. We found that in both models the input data performs very well by capturing the phase transitions accurately. This highlights the versatility of our input data in identifying conventional phase transitions, both continuous and discontinuous, as well as topological phase transitions. Moreover, we employed the spin-opstring to perform a regression task predicting the superfluid density in the $2D$ $t_{2}-V_{1}$ model. We have further compared the performance of the spin-opstring with the full simulation cell data. Our results show that as system size increases, the combined memory and computational costs make the full simulation cell impractical.
@@
To explore the broader applicability of the spin-opstring, we conducted transfer learning experiments across different physical models. First, a model trained on the $t_{2}-V_{1}$ system was successfully applied to predict phase transitions in the $t_{1}-V_{1}$ model which features a similar superfluid–checkerboard solid transition but originates from nearest-neighbor (NN) hopping rather than next-nearest-neighbor (NNN) hopping. Despite this difference in origin, the model can very well apply the learned knowledge from the $t_{2}-V_{1}$ model to identify the phase boundary in the new system. In a second case, a ML model trained on the non-interacting SSH model accurately predicted phase transitions in the interacting SSH model with nearest-neighbor repulsion, up to some interaction strength. These results illustrate the strong potential of spin-opstring data in facilitating knowledge transfer between related models. We also evaluated generalization across system sizes, by training a ML network on the SSH model in a smaller lattice and testing it on a larger system. The model correctly identified the topological-to-trivial phase transition, confirming that spin-opstring data supports generalization to larger systems.
@@
Overall, our results establish that the combined spin-opstring can successfully be applied across a wide range of ML tasks. For completeness we also addressed the question whether either the spin-state or the operator string could serve as stand-alone inputs. We considered examples to demonstrate that the unified spin-opstring representation leads to a more reliable model performance than using either component individually.
@@
Finally, we employed two leading interpretability techniques, Layer-wise Relevance Propagation (LRP) and SHapley Additive exPlanations (SHAP), to gain insight into the model’s decision-making processes. We devised a strategy to make the high-dimensional outputs of LRP and SHAP interpretable by linking the feature scores to specific physical operators. Our analysis revealed that while LRP disregards identity operators (as they are encoded as zeros in the spin-opstring), SHAP is able to recognize their contextual importance. This highlights the advantage of SHAP in scenarios where every component of the input, including identity operators, contribute to the model’s predictions. Across classification, regression, and transfer learning tasks, our interpretability study consistently showed that the models learned physically meaningful features from the input data. This supports the conclusion that the spin-opstring is not only effective but also interpretable for ML-based analysis of quantum systems.

### v3

@@ VIII Conclusions
In conclusion, we have proposed to use the “spin-opstring”, extracted from Stochastic Series Expansion QMC as a input data type for machine learning applications. The spin-opstring is a compact representation of the QMC simulation cell that retains all quantum information of the system while being memory-efficient. It consists of the initial state of the simulation cell followed by an operator string that describes the state’s evolution through imaginary time, ultimately bringing it back to its original configuration. To demonstrate the applicability of spin-opstring as ML input, we consider two different models : the 2D $t_{2}-V_{1}$ model and the 1D SSH model. We find that trained ML models are able to successfully predict conventional and topological phase transitions. In a regression setting, ML models were also able to correctly predict the superfluid density in the 2D $t_{2}-V_{1}$ model. We have further compared the performance of the spin-opstring with the full simulation cell data. Our results show that as system size increases, the combined memory and computational costs make the full simulation cell impractical.
@@
We showcase broader applicability of spin-opstring with successful transfer learning experiments. First, we show that a ML network trained on the non-interacting SSH model accurately predicted phase transitions in its interacting version up to moderate interaction strength. In another example, a model trained on the $t_{2}-V_{1}$ system correctly predicted a similar phase transition in the $t_{1}-V_{1}$ system. These results illustrate the strong potential of spin-opstring data in facilitating knowledge transfer between related models. Furthermore, by training a ML network on a smaller lattice and predicting transitions on a larger system, we confirm that spin-opstring supports generalization across system sizes.
@@
Overall, our results establish that spin-opstring can successfully be applied across a wide range of ML tasks. However, this raises the question of whether it can surpass conventional inputs based on spin configurations. In a HCB system, dubbed as the 2D WTI model, we demonstrate the superiority of spin-opstring over spin-state as ML input in accurately predicting a quantum phase transition. We also present a case where using the opstring alone as input causes training issues. These findings demonstrate that the unified spin-opstring representation leads to more reliable model performance.
@@
Finally, we employed two leading interpretability techniques, Layer-wise Relevance Propagation (LRP) and SHapley Additive exPlanations (SHAP), to gain insight into the model’s decision-making processes. Leveraging the internal structure of spin-opstring, we devised a strategy to make the high-dimensional outputs of LRP and SHAP interpretable by linking the feature scores to specific physical operators. Our analysis revealed that while LRP disregards identity operators (as they are encoded as zeros in the spin-opstring), SHAP is able to recognize their contextual importance. This highlights the advantage of SHAP in scenarios where every component of the input, including identity operators, contribute to the model’s predictions. Across classification, regression, and transfer learning tasks, our interpretability study consistently showed that the models learned physically meaningful features from the input data. This supports the conclusion that the spin-opstring is not only effective but also interpretable for ML-based analysis of quantum systems.
@@ Appendix A Effect of Batch Normalization
Appendix B Domain adaption from $t_{2}-V_{1}$ to $t_{1}-V_{1}$
To explore the application of transfer learning we consider another example; the well-known $t_{1}-V_{1}$ model on the periodic square lattice, governed by the Hamiltonian,
$H=-t_{1}\sum_{\langle i,j\rangle}(b^{\dagger}_{i}b_{j}+\textrm{h.c.})+V_{1}\sum_{\langle i,j\rangle}n_{i}n_{j}-\mu\sum_{i}n_{i}.$ (14)
Appendix C Details of architecture used in generalization across system sizes
A possible way to perform the training for this task would be to equalize the spin-opstring lengths of the larger and smaller systems through padding. However, this method is not compatible with arbitrary length data during inference and it also has an additional computational overhead. Instead, to handle the variable length data, we have incorporated the Global Pooling Layer in 1D CNN architecture. In particular for the 1D SSH model we keep the same network structure as before but replace the flattening layer with a global average pooling layer. This layer reduces output of each CNN filter to a single value, which is an average over all its entries, thus effectively converting inputs of varying lengths into fixed-size output vectors. The subsequent FC layers always receive inputs of consistent size, allowing the model to generalize across different system sizes without needing padding or truncation.
Appendix D Interpretability analysis using Layer-wise Relevance Propagation
Figure 22: LRP analysis of the cS-cSS-MI transition as a function of chemical potential $\mu$ with repulsion fixed at $V_{1}=4.0$. Top: LRP sums for the spin-state (scaled for clarity) and the operator string. Bottom: LRP contributions from diagonal and off-diagonal operators; identity operators are omitted due to zero relevance. The spin-state and off-diagonal sums are scaled for better visibility. In both plots, the predicted transition points are indicated by dashed vertical lines.
Layer-wise Relevance Propagation (LRP) is a powerful post hoc interpretability technique that decomposes the prediction of the model by propagating relevance backward from each output neuron to the input features [29]. With the application of a set of conservation-based propagation rules at each layer, LRP yields a relevance map over the input data.
$R_{j}=\sum_{k}\frac{z_{jk}}{\epsilon+\sum_{j}z_{jk}}R_{k}\,$, (15)
where $R_{j}$ corresponds to the score for a neuron $j$ in the inner layer and $R_{k}$ for a neuron $k$ in the outer layer. The coefficients $z_{jk}$ represent the model weights and biases present in between the two layers. The core idea of the LRP-$\epsilon$ rule involves adding a small positive constant $\epsilon$ to the denominator, the role of which is to offset some relevance from weak or contradictory neuron activations. Thus it helps to retain the most salient relevances by suppressing noise and avoids numerical instabilities arising from vanishing activations. We have used $\epsilon=10^{-3}$ for the LRP-$\epsilon$ rule and performed our LRP analysis using the iNNvestigate library [43] which allows integration with Keras.
We analyse the phase classification networks in the 2D $t_{2}-V_{1}$ model using LRP. Similar to SHAP, we use ML classifiers without the final softmax activation for the analysis. The interpretability results obtained from LRP are mostly consistent with SHAP with one major difference, LRP assigns zero relevance to identity operators and is therefore unable to ascertain their role. Consequently, we omit identity operators in the LRP figures below.
In Fig. 21, we show the total LRP sums for spin-state and opstring (top), and the individual operator types (bottom) across the SF–cS transition. Consistent with the SHAP analysis, the spin-state is far less important for this classification than the opstring. The spin-state’s relevance drops to zero in the SF phase, likely due to the absence of any spin-order. In the cS phase, however, its positive and negative relevances correspond to the two degenerate spin configurations. The LRP scores for diagonal and off-diagonal operators show a qualitatively similar trend to SHAP results. However, as mentioned earlier, LRP is unable to identify the role of identity operators in the cS phase.
Next, for the cS-cSS-MI phase transition, we show the LRP relevance scores for the spin-opstring components in Fig. 4. In the cSS and MI phases, the score trends for spin-state, opstring and different operators are similar to those from the SHAP analysis. In the cS phase, however, a discrepancy arises, the total LRP sum for the spin-opstring is negative, which contradicts the positive logit value from the model’s output. This result appears to be a significant violation of the LRP conservation property.
The source of this discrepancy becomes clear when we compare the LRP analysis with the SHAP results (Fig. 16). In the cS phase, both methods assign negative scores to diagonal operators, while off-diagonal ones remain zero. The total SHAP sum, however, becomes positive due to a large positive contribution from identity operators. Since LRP assigns zero relevance to these operators, its total score fails to match the positive logit value. This contradiction originates from a fundamental difference in how the two methods attribute relevance. In the spin-opstring data, identity operators are encoded as zero. Since LRP propagates relevance based on neuron activations and weights, it tends to assign zero relevance to features that are numerically zero in the input [44], regardless of their contextual role. In contrast, SHAP assigns non-zero relevance to identity operators by measuring how its presence or absence influences the model’s prediction.